### v1

# Seeing the Trees for the Forest: Rethinking Weakly-Supervised Medical Visual Grounding

Abstract
Visual grounding (VG) is the capability to localize specific regions in an image associated with a particular text description. In medical imaging, VG enhances interpretability and cross-modal alignment of multimodal models by highlighting relevant pathological features corresponding to textual descriptions. Current vision-language models struggle to associate textual descriptions with disease regions due to inefficient attention mechanisms and a lack of fine-grained token representations. In this paper, we empirically demonstrate two key observations. First, current VLMs assign high norms to background tokens, diverting the model’s attention from regions of disease. Second, the global tokens used for cross-modal learning are not representative of local disease tokens. This hampers identifying correlations between the text and disease tokens. To address this, we introduce simple, yet effective Disease-Aware Prompting (DAP) process, which uses the explainability map of a VLM to identify the appropriate image features. This simple strategy amplifies disease-relevant regions while suppressing background interference. Without any additional pixel-level annotations, DAP improves visual grounding accuracy by 20.74% compared to state-of-the-art methods across three major chest X-ray datasets.
1 Introduction
Visual grounding (VG) is the task of localizing the specific region in an image that corresponds to a given text description. VG is a prominent research area at the vision-language intersection [14, 17, 31], enabling applications like visual question answering  [23, 52], human-robot interaction [47, 51], and data annotation [24, 56]. The ability of linking textual mentioned diseases to specific pathological image regions enhances interpretability [53], aiding radiologists in assessing model decisions. Additionally, VG integrates effectively with large language models (LLMs) to enhance multimodal understanding, e.g., refining 3D object localization through interleaved reasoning and grounding [67], summarizing long text with LLM to improve VG [68], and introducing iterative vision-language interaction with novel tokens for object detection (VG) and text generation [33]. In summary, VG enables transparency, assists localization, and enhances multi-modal interactions on the patch level of the emerging medical multi-modal models [39, 11, 19].
Many methods [10, 24, 39, 21] perform supervised visual grounding on medical imaging. They rely on text-and-region correspondence annotations to train a cross-modal fusion module to ground text to specific disease regions. However, supervised approaches require dense annotations at the pixel level, which is expensive in the medical settings. Another work [5] proposes weakly-supervised VG for a pneumonia by using a pre-trained object detector.
Figure 1: Left: Average cosine similarity between disease patch tokens and the corresponding disease texts of CheXzero [50] on MS-CXR [6] dataset. Without disease-aware prompting, the current VLM shows undifferentiated alignment scores between disease and non-disease patches with text. Right: Local feature norms, and visual grounding of CheXzero [50], Self-EQ [22], and our disease-aware prompted model.
Recently, weakly-supervised visual grounding methods [29, 38, 9, 35, 36, 43, 49, 37] have leveraged strong vision-language models (VLMs) such as CLIP [42], as a backbone, and achieve a solid visual grounding performance without using text-to-region correspondence labels. Most methods obtain localization pseudo-labels using explainability methods [34, 8] applied to the vision-language models. Alternatively, mask-driven methods train a segmentation decoder to predict the foreground and background masks of CLIP [45, 44], or predict the masks used to mix two images [2]. Overall, current methods get the patch-level features from the pre-trained VLM backbone on the images then aim to improve local alignment. Unlike general images that typically contain multiple distinct objects , medical images typically contain more ’stuff’[7], and thus present a unique challenge: critical disease indicators often occupy only a small proportion of the image, while anatomical backgrounds dominate visual features. The trees are lost in the forest. As such, applied to medical imaging, current VLMs [50, 41, 4, 6] yield sub-optimal local features, which are less disease-focused.
In this work, we set out to analyze this limitation in current medical VLMs, identify whether it hinders visual grounding methods, and if so to propose improvements. Our analysis shows that: 1) We observe that the $L_{2}$ norm of the token embeddings are high in background non-disease regions , as shown in Fig. 1. As the norm reflects token vector magnitude and representation strength, high background norms can divert attention from disease-relevant areas, impairing medical visual grounding. 2) As disease regions are typically small, global image tokens are more strongly aligned with the major non-disease image patch tokens, and less representative of the fine-grained disease tokens, as shown in Fig. 8. As such, here is low alignment between disease image tokens and the global text and image tokens, which impairs visual grounding.
Figure 2: In current VLMs, text [CLS] ( $\bullet$ ) are pulled close to a global image token [IMG] ( $\bullet$ ) through contrastive loss, and thus they use the global token as a proxy to link with disease patch tokens ( $\bullet$ ). Left: w/o prompt, the global image token and the disease patch of the original input on the left show weak alignment, while our proposed disease-aware prompting enhances contrast for disease regions, aligning disease patches with both [IMG] and text [CLS]. Right: The original and prompted feature space for the images having disease ‘atelectasis’ on MS-CXR [6] dataset. The disease tokens are pulled towards the [IMG] upon prompted.
The two problems identified above illustrate the influence of non-disease background regions on the medical visual grounding challenge. To remedy their impact, we propose a simple-yet-effective disease-aware feature prompting approach, as illustrated in Fig. 2. We first extract the explainability map [8] of the pre-trained VLM for the given textual disease description. We then use this disease-aware map and prompt the vision models on the feature space, thus suppressing features of non-disease tokens and amplifying disease-related tokens. The model then extracts a global token, i.e. [IMG] token from this disease-focused feature map, bringing disease patches closer to the global image [IMG] token and improving its alignment with the text [CLS] token. This design is inspired by how radiologists modify the contrast to make the disease regions more salient and localize them. As shown in Fig. 1 (Right), our prompting strategy effectively assigns high norm values to disease regions, producing accurate visual grounding predictions. Without using any dense annotation, our method achieves superior performance compared to several state-of-the-art methods on multiple datasets.
Our main contributions are summarized as follows:
  • We show that medical VLMs have high norm values in non-disease background regions, which misdirects the visual grounding of fine-grained disease regions.
  • We further show that current VLMs suffer from intra-modal misalignment, where the global image token fails to capture the disease information from the local disease tokens. This weakens local patch-to-text alignment when training via global image-text contrastive losses.
  • We introduce a simple-yet-effective disease-aware feature prompting for weakly-supervised medical visual grounding. Our technique suppresses the influence of non-disease local tokens to concentrate the model attention in relevant disease regions.
  • Our proposed method achieves superior performance compared to several state-of-the-art methods for weakly-supervised medical visual grounding across three datasets. This paper is the first work, establishing a comprehensive medical weakly-supervised visual grounding benchmark on three datasets.
2 Related Works
Weakly-Supervised Visual Grounding has the same goal as visual grounding but is achieved using only image-level annotations. Prior methods [55, 37, 20, 54, 13] for weakly-supervised visual grounding rely on pre-trained object detectors for Region of Interest (ROI) localization. They typically project text and ROIs into a joint visual-textual embedding space, transforming the grounding task into a retrieval task. In contrast, detector-free methods [45, 44, 2] perform dense localization for a given query phrase without relying on a pre-trained masks proposal model. There are two main approaches in detector-free methods. First, many works [45, 44, 2, 34] generate pseudo-labels, which are expected to approximate the ground truth segmentation, and train models on visual grounding objective. G [45] and g++ [44] employ layer-wise relevance propagation [8] to integrate relevancy map into the cross-modal fusion. Second, other methods [26, 59, 63] optimize some auxiliary tasks on the weakly supervised data. SelfEQ [22] tune model explanation, thus having no detector or pseudo-label generation. Although visual grounding has been well explored for natural images, studies of its applicability in the medical domain is not abundant. Recently, few works [10, 24] perform supervised medical visual grounding that requires pixel-level annotations, which is expensive in the medical domain.
Vision-language models have garnered significant interest recently, largely since the success of CLIP [42]. Other notable VLMs include ALIGN [27], Florence [62], and Flamingo [1] have pushed forward research in this area. Building on these advances, several medical VLMs have emerged. These models learn the cross-modal information from the vision and language modalities such as chest radiographs and associated reports. CheXzero [50], MedCLIP [57], and BiomedCLIP [64] align images with raw medical reports with CLIP’s contrastive learning objective, while BioVIL [6] and BioVIL-T [4] enhance the text encoding process. MedKLIP [58] and MAVL [41] leverage fine-grained report structures from RadGraph [25] to further improve performance. MedIM [60] introduces masked image modelling in medical VLMs. Given the fine-grained nature of medical images, where diseases often occupy small areas, current medical VLMs struggle to associate thefine-grained pathological regions with textual descriptions.
Prompt tuning is a strategy to adapt large-scale pre-trained models to new tasks. CoOp [66] and CoCoOp [65] prepend the text tokens with learnable parameters to adapt CLIP for many recognition tasks. The textual prompt tuning is widely used in open-vocabulary object detection [18] and semantic segmentation  [61]. Several works [34, 28, 3] apply visual prompt tuning on the pixel space [3, 61] and feature space [28], to adapt a frozen pre-trained model. Previous works adopt learnable, and randomly initialized prompts for model adapting. In contrast, we adopt informative prompting, which enforces rich disease-aware prompts to guide medical visual grounding.
3 Proposed Method
Problem formulation. Given an input image $x_{\text{v}}\in\mathbb{R}^{C\times H\times W}$ and a text query $x_{\text{t}}$ specifying an object in $x_{\text{v}}$, weakly supervised visual grounding aims to produce a segmentation map $y\in\mathbb{R}^{H\times W}$ for the referenced object, without dense segmentation labels. The visual grounding model leverages a VLM with an image encoder $F_{\text{v}}$ and text encoder $F_{\text{t}}$, followed by a pixel decoder $D$. $F_{\text{v}}$ provides a sequence of local visual token features $V=\{v_{1},\ldots,v_{n}\}$ and the global image token [IMG] from $x_{\text{v}}$ , while $F_{\text{t}}$ yields text [CLS] token from $x_{\text{t}}$:
$\{[\texttt{IMG}],v_{1},v_{2},\dots,v_{n}\}=F_{\text{v}}(x_{\text{v}}),\quad\texttt{[CLS]}=F_{\text{t}}(x_{\text{t}}).$ (1)
3.1 Problem Analysis
Current VLMs [50, 4, 41] is efficient at identifying what diseases are present, i.e. image classification. However, it struggles in visual grounding, determining where to align text descriptions on the medical images. Here, we refer to the target disease regions as foreground (FG), and non-disease regions as background (BG). Fig. 1 (Left) shows the cosine similarity distributions of text tokens with the FG and BG tokens when using the state-of-the-art CheXzero [50]. It shows that BG tokens have a higher alignment with the text.
VLMs [50, 6, 4] are mostly trained using only the global image token [IMG], i.e., extracting features from the entire image and aligning with the text [CLS]. However, medical images often feature multiple diseases with small regions, making it difficult to correctly assign text to the corresponding regions. This brings challenges in fine-grained visual grounding. Fig. 3 shows the visual grounding performance of the VL model on cases with multiple and small diseases. The Dice drops as the disease areas reduce, and the number of present diseases increases.
Figure 3: Performance analysis of BioVIL [6] on MS-CXR dataset as a function of (a) disease region size and (b) disease count per image. Left: Dice scores stratified by disease area, showing degraded performance for smaller regions. Right: Performance impact of concurrent diseases, demonstrating declining accuracy with increasing disease count. Disease areas are categorized into five groups by size percentiles.
Figure 4: Comparative analysis of token norm distribution (Norm) and visual grounding (VG) performance across methods with representative cases show common chest pathologies. Left to right: Ground truth (GT), CheXzero [50] (foundational VLM), SelfEQ [22] (weakly-supervised baseline), and our disease-aware prompting approach. Our method produces more focused norm activations that better correspond to disease regions, leading to more precise grounding results. Colors indicate activation strength from low (blue) to high (red). Text descriptions are abbreviated.
High norm on BG tokens, hampering the visual grounding capability. The $L_{2}$ norm of patch tokens $\|v\|_{2}$ indicates the activations of the local token $v$. Fig. 4 shows the norm of patch tokens of CheXzero. Tokens in BG regions are frequently activated with high values. Unlike natural images, medical images are highly canonical, with significant redundancy in BG areas. Hence, high-norm tokens often occur in these healthy BG regions rather than in the disease-associated FG areas. We then show that the high activation of BG tokens hampers the visual grounding. Fig. 4 further compares the norm map and the visual grounding of CheXzero [50]. The model mistakenly grounds on background tokens with high norms.
To quantify how spurious norm activation in the background (BG) regions affects visual grounding (VG), we calculate the $L_{2}$ norm of the tokens to form a norm map. We then evaluate the correlation between these norm maps and VG predictions. Fig. 5 shows two findings. First, VLMs show high norm activations in the background regions, evident by the high Dice score between the response map and the BG regions. Second, the Dice score between VG outputs on the norm maps is significantly higher than the Dice with the actual ground truth (GT). This shows that the models tend to ground on the highly activated regions of the norm map. These findings suggest that high norm activations in BG regions can divert the model away from disease areas, weakening VG accuracy.
Figure 5: Dice score (%) between (i) Norm map and the background regions (BG); (ii) the visual grounding output (VG) and the norm map; and VG output and ground-truth (GT). Analysis of recent vision-language pre-training models [50, 6] and a recent weakly-supervised visual grounding method [22].
Global image [IMG] token is non-representative of the local foreground disease tokens. Current VLMs mostly operate on the global image [IMG] when aligning with the text token. Are they representative of local patch tokens $v_{i}$? This study shows a negative finding. Fig.6 investigates the intra-modal alignment between the global [IMG] token and foreground (FG) and background (BG) tokens using cosine similarity. In the top row, the [IMG] token aligns more closely with BG tokens than with FG tokens. This is further substantiated by the average cosine similarity between the global [IMG] token and individual patches (see Fig. 7).
Figure 6: Intra-modal alignment. Distribution of cosine similarity scores between the global image [IMG] token and the image patch tokens, for original input and prompted input. Training with the disease-aware prompts boosts the intra-modal alignment between the global token [IMG] and disease patch tokens, indicated by the rightward skew in the distribution of cosine similarity between [IMG] and FG patches.
Figure 7: Average cosine similarity between the global image [IMG] token and the image patch tokens. Left: Original CheXzero model shows higher similarity to background without prompt than disease regions. Right: After prompting, the relationship inverts with stronger disease region alignment.
Current VLMs use the global image [IMG] token as an intermediary between the text [CLS] token and the local image patches. Hence, if the global image [IMG] token is non-representative of fine-grained local disease information, the alignment between the text [CLS] and FG tokens is suboptimal. To validate, we show the feature distribution of global image [IMG], text [CLS], BG, and FG tokens in Fig. 8. Due to the global cross-modal contrastive objective, the image and text [CLS] tokens are positioned closely to each other, yet both remain distant from FG disease tokens.
Figure 8: Feature space visualization for six common chest conditions without (Top) and with our disease-focused prompting(Bottom). The global image token ( $\bullet$ ) is close to the text token ( $\bullet$ ). However, the disease foreground patch ( $\bullet$ ) tokens are far from the global image token, thus being far from the text token. In contrast, our feature space shows higher alignment between local and global tokens, and local tokens and the text token.
Figure 9: Model architecture. We use the disease-aware prompt to align the global image token [IMG] and [CLS] for the Disease-aware global contrastive loss, and to select FG/BG patch tokens for the Local Contrastive Loss. During inference, only the original image patch token and text [CLS] are used.
3.2 Disease-aware Prompting for Weakly-Supervised Visual Grounding
The study identifies two key issues: (1) weak alignment between the local tokens $V$ and global [CLS] token, which fails to represent disease-relevant regions, and (2) high-norm background (BG) tokens that undesirably influence the visual grounding. As such, we propose a simple-yet-effective disease-aware feature prompting approach to enhance the influence of disease regions and reduce the impacts of background tokens. From the input image-text pair, we first extract the explainability map of the pretrained VLM. Our framework uses this explainability map to prompt the image features, enhancing disease token contributions to the global [IMG], suppressing irrelevant BG’s influence. Built upon this disease-aware prompting map, we adopt a hierarchical training objective with global and local contrastive losses as illustrated in Fig. 9.
Disease-aware feature prompting. To prioritize disease regions in the model’s attention, we apply prompting to reweight the visual token patches so that disease FG regions have stronger impact on the global image [IMG] tokens within the Vision Transformer (ViT) via self-attention. Employing the VLM interpretability approach [8], we generate a 2D Disease-aware Prompt $\Phi$ as follows:
$\Phi=\mathcal{R}\Big{(}V^{l},\,\mathcal{M}\big{(}F_{\text{v}}(x_{\text{v}}),\,F_{\text{t}}(x_{\text{t}})\big{)}\Big{)},$ (2)
where $V^{l}=\{v^{l}_{1},\ldots,v^{l}_{n}\}$ denotes the intermediate visual tokens features at the layer $l$ to interpret, $\mathcal{M}$ represents the matching function between the multi-modal outputs of the image and text encoders, and $\mathcal{R}$ applies the interpretability approach [8]. Although imperfect, $\Phi$ can effectively highlight disease-relevant FG regions indicated by the text $x_{\text{t}}$. We treat $\Phi$ as the disease-specific importance map for $x_{\text{t}}$, using it to weight the visual tokens features, thereby amplify the FG disease signals while suppressing BG noise:
$\hat{V}^{l}=\Phi\cdot V^{l}.$ (3)
Accordingly, the disease-aware prompted visual output at the final layer is:
[…]
The visual grounding output is produced by the pixel decoder, where we employed cross-attention to fuse the text and vision features:
$\hat{y}=D\big{(}\text{CrossAttn}(V^{L},\texttt{[CLS]}^{L})\big{)}.$ (5)
Figure 10: Local contrastive loss. The FG and BG tokens are selected using the Disease-aware prompt $\Phi$.
Loss objectives. We introduce hierarchical training objective with global and local losses to maximize granularity. First, we introduce the disease-aware global contrastive loss. After concentrating local disease information into the global image [IMG] token via disease-aware prompting, we position disease-related local tokens close to the global image [IMG]. As such, using a global contrastive loss $\mathcal{L}_{\text{glb}}$, we bring the text [CLS] closer to the image [IMG], which serves as a proxy to align disease-related local image tokens with the text [CLS]. Given the batch of image-text pairs $\{(x_{\text{i}},x_{\text{t}})\}^{n}_{i=1}$, the cross-modal global contrastive loss is:
$\mathcal{L}_{\text{glb}}=-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\exp\left[\cos\left(\texttt{[IMG]}^{L}_{\Phi_{i}},\texttt{[CLS]}_{i}\right)\right]}{\underset{j\in\{x_{t}^{-}\}}{\sum}\exp\left[\cos\left(\texttt{[IMG]}^{L}_{\Phi_{i}},\texttt{[CLS]}_{j}\right)\right]},$ (6)
where $\{x_{t}^{-}\}$ indicates the set of negative text descriptions of $x_{i}$ in the batch.
Second, we employ the local contrastive loss $\mathcal{L}_{\textit{lcl}}$ to align textual descriptions with disease regions in the image, as shown in Fig. 10. We threshold the disease-aware prompting mask $\Phi$ to select FG tokens $\text{FG}(V)=\{v^{L}\}_{\text{FG}}$ and draw them closer to the text [CLS], while pushing BG tokens $\text{BG}(V)=\{v^{L}\}_{\text{BG}}$ away. The local contrastive loss for each image-text pair is defined as:
$\mathcal{L}_{\text{lcl}}=-\frac{1}{|\text{FG}_{i}|}\sum_{v_{\text{fg}}\in\text{FG}_{i}}\log\frac{\exp\left[\cos\left(v_{\text{fg}},\texttt{[CLS]}_{i}\right)\right]}{\sum\limits_{v_{\text{bg}}\in\text{BG}_{i}}\exp\left[\cos\left(v_{\text{bg}},\texttt{[CLS]}_{i}\right)\right]}.$
Figure 11: Qualitative results of recent VLMs, and weakly-supervised VG methods on MS-CXR[6] , RSNA [46] and Covid19 [15].
Finally, we learn the pixel decoder with $\Phi$ as pseudo-label, using Dice loss $\mathcal{L}_{\text{seg}}$. The total loss objective is:
$\mathcal{L}=\mathcal{L}_{\text{glb}}+\mathcal{L}_{\text{lcl}}+\mathcal{L}_{\text{seg}}$ (7)
Note that, the extracted disease-aware prompt $\Phi$ is only needed during training to resolve the local misalignment of VLM. Our disease-aware contrastive loss $\mathcal{L}_{\text{glb}}$ pulls the text [CLS] close to the more disease-representative global token [IMG], thus improving the local alignment between text and local disease tokens $V$. Hence, during inference, our model can ground disease patches on the original image without relying on the pre-extracted explainability map. Using this map to narrow down the pathological search space, our DAP enables the model to identify subtle and fine-grained disease features, thus further improving the disease visual grounding.
Discussion. While a recent supervised visual grounding method, ZSeg [61], also proposes to mask the input image, our disease-aware prompting differs. ZSeg masks in the pixel space; in contrast, we propose to prompt in the feature space to address the followings. First, pixel hard-masking can erase relevant context, while we multiply $\Phi$ with token features, preserves both local and global information. Second, pixel masking is insufficient to suppress high background norms as it only operates at the image level. Our method effectively reduces high background norm compared to pixel masking as shown in Fig. 12.
Table 1: Weakly-supervised visual grounding benchmark of different methods on MS-CXR [6] and RSNA [46] on CNR as our main metric, and other standard visual grounding metrics.
Dataset MS-CXR RSNA Pneumonia
[…]
CheXzero [50] Nat. BE 22 0.675 0.156 0.163 0.089 0.814 0.468 0.202 0.120
BioViL [6] ECCV22 0.941 0.326 0.238 0.216 1.096 0.472 0.339 0.252
BioViL-T [4] CVPR23 1.164 0.416 0.320 0.218 0.805 0.584 0.295 0.192
Trans-MM [8] ICCV21 1.122 0.406 0.336 0.244 1.293 0.526 0.421 0.284
[45] NIPS22 0.672 0.304 0.290 0.186 1.461 0.284 0.450 0.315
g++ [44] CVPR23 0.569 0.258 0.268 0.152 1.350 0.467 0.445 0.311
VPT [34] ICASSP24 0.588 0.323 0.255 0.162 1.173 0.612 0.468 0.327
Self-EQ [22] CVPR24 1.082 0.406 0.327 0.229 1.075 0.637 0.370 0.241
DAP Proposed 1.254 0.457 0.352 0.245 1.630 0.747 0.474 0.328
4 Experimental settings
Datasets. MIMIC-CXR v2 chest radiograph data [30] comprises 227,835 imaging studies from 65,379 patients with image-level labels. We include only frontal-view scans, utilizing in a train/validation set of 146.7k/22.2k samples. Evaluation was done using MS-CXR [6], RSNA Pneumonia [46] and COVID Rural [15] datasets. MS-CXR [6] is a subset of MIMIC, provides bounding boxes for 1162 images on 8 diseases. RSNA Pneumonia [46] includes 26k frontal chest X-rays with pneumonia masks, provided by the Radiological Society of North America. COVID Rural [15] contains over 200 annotated chest X-rays for COVID-19 segmentation. We use GPT4-o [40] to generate 50 unique text descriptions per disease label, which are reviewed by board-certified radiologists to create a refined text description database. During training, models are prompted with a random sample from its respective disease descriptions set.
Comparison approaches. We benchmark two groups. The first group is the medical VL pre-training models that are trained on a large-scale image-report pairs, including BioViL [6], BioViL-T [4], and CheXzero [50]. The second group includes state-of-the-art weakly visual grounding models, which are trained via the pseudo-labels generated from the pre-trained VLMs. These include Self-EQ [22], g [45], g++ [44], Transformer-MM [8], and VPT [34]. Since medical images do not have a strong and universal disease object detectors, we only benchmark against detector-free VG methods, and exclude detector-based methods [29, 37].
Metrics. Following recent medical visual grounding methods [6, 4], we adopt the contrast-to-noise ratio (CNR) as the main metric, which measures the difference between FG and BG activations. Besides, we adopt standard VG metrics, including pointing game (PG), Dice and IoU scores.
Implementation Details. Our experiments are conducted on 4 NVIDIA RTX A100 GPUs. Input images are resized to 224$\times$224 and augmented with RandAugment [12]. We set up an Adam optimizer [16] with a learning rate of 0.008 and a batch size of 512 across all experiments. We cap the maximum length of each text to 77 as suggested by [50]. Our hyperparameter values and schedules were determined empirically on a small validation subset.
5 Results and Discussions
Comparisons with state-of-the-art methods. Table 1 shows the results of recent methods on the two datasets. Our method demonstrates a notable improvement over the vision-language model BioViL, achieving a CNR boost of 7.73% and 48.72% respectively on MS-CXR and RSNA datasets. When compared with Self-EQ [22], a recent advancement in weakly-supervised visual grounding, our approach significantly outperforms it by 15.89% in CNR and 9.36% in the pointing game on MS-CXR. Furthermore, we surpass g++ [44] by 20.74% in CNR on the RSNA dataset.
Table 2 presents the comparisons on the Covid-19 segmentation dataset. We significantly outperforms the recent weakly-visual grounding methods, VPT, and Self-EQ by 17.7% and 33.26% in terms of PG. Fig. 11 shows that existing visual grounding methods tends to over-segment the disease regions. These results underscore that existing weakly-supervised visual grounding methods fall short on medical imaging tasks, mainly due to the limited disease-focused local features within the VLM backbone. In contrast, ours show crisp localization on the fine-grained disease regions.
Table 2: Weakly-supervised visual grounding benchmark of different methods on Covid19 [15].
[…]
CheXzero [50] Nat. BE 22 0.923 0.235 0.165 0.097
BioViL [6] ECCV22 0.623 0.438 0.177 0.201
BioViL-T [4] CVPR23 0.739 0.469 0.297 0.220
Trans-MM [8] ICCV21 0.955 0.406 0.325 0.223
[45] NIPS22 0.610 0.282 0.246 0.157
g++ [44] CVPR23 0.702 0.469 0.311 0.205
VPT [34] ICASSP24 0.739 0.531 0.332 0.220
Self-EQ [22] CVPR24 0.659 0.469 0.304 0.190
DAP Proposed 1.018 0.625 0.391 0.267
Figure 12: Token norm and VG analysis of prompt tuning methods: VPT [28], Pixel mask [61] versus our DAP.
Ablation Study. Norm and visual grounding of prompt tuning methods. We evaluate the token norms of visual prompts in VPT [34] and pixel prompts in ZSeg [61].VPT only learns the prompt and does not overcome the problem of high norms in BG regions. Pixel prompting reduces BG noise at the pixel level, but its effect weakens in deeper layers. Our approach, which prompts directly in the feature space, provides consistent background suppression, with improved qualitative results demonstrated in Fig. 12.
Comparisons of different prompting. We compare VG performance of different prompting techniques, as shown in Table 3. Notably, we compares ours with pixel prompting adopted in ZSeg [61], which applies object binary masks on the image pixels. Ours significantly improves upon the pixel masking prompting by 32% in CNR metrics. Ours explicitly suppresses the non-disease regions, thus optimally alleviating the influence of background regions. Secondly, our disease-aware prompting outperforms learnable textual prompts (e.g., CoCoOp [66]), visual prompts [28], and multi-modal prompts [32]. This shows that informative disease-aware prompting is more suitable for fine-grained medical visual grounding.
Table 3: Comparisons of different prompt tuning techniques applied on weakly-supervised visual grounding on the MS-CXR dataset [6].
[…]
Red circle [48] 0.866 0.307 0.261 0.177
Pixel mask [61] 0.949 0.361 0.306 0.199
CoOp [66] 0.986 0.364 0.311 0.206
CoCoOp [65] 1.042 0.418 0.333 0.226
Visual prompt [28] 1.129 0.418 0.338 0.231
Multi-modal [32] 1.101 0.414 0.345 0.233
Disease-aware (Ours) 1.254 0.457 0.352 0.245
Ablation study of the proposed components. Table 4 shows the performance when ablating different components. First, adding disease-aware prompting into the global contrastive loss significantly improves the Dice score by 24.01%. This shows that using VLM with only global contrastive loss can lead to the low local alignment, hindering the visual grounding. Second, integrating DAP and global contrastive loss yields comparable performance with the local contrastive loss. Without fine-grained local objective, our disease-aware prompting helps pull the local patches closer to [CLS] only using the global token [IMG] as a proxy. Applying all components yield the highest performance, showing the complementary of disease-aware prompting and global-to-local objectives for fine-grained alignment.
Table 4: Ablation study of the disease-aware prompting (DAP), the global contrastive loss, and the local contrastive loss on the MS-CXR dataset.
[…]
0.878 0.376 0.279 0.153
1.061 0.441 0.346 0.237
1.064 0.434 0.343 0.238
1.254 0.457 0.352 0.245
6 Conclusion
This paper provides a diagnostic view of local alignment of medical vision-language models, hampering the weakly-supervised visual grounding. We show that (i) these models assign high norm values to non-disease regions, misdirecting attention, and (ii) intra-modal misalignment between global and local tokens weakens disease-specific grounding. As such, we propose a simple-yet-effective disease-aware prompting (DAP) technique that strengthens disease-region representations while suppressing background interference. This paper constructs a comprehensive weakly-supervised medical visual grounding on three datasets to catalyse the research. On this new benchmark, ours outperforms state-of-the-art methods by up to 20.74% across 3 datasets.